What mass of Lithium Oxide is formed when 1.0g of Lithium is burned in 1.3g of Oxygen


1
Expert's answer
2021-09-19T00:03:54-0400

Mole ratio = 1:1

Molar mass of Lithium = 6.941

1.0/6.941 = 0.1441mol

Molar mass of Oxygen = 15.999

1.3/15.999 = 0.0813

Molar mass of Lithium oxide = 29.88

0.0813 + 0.1441 = 0.2254× 29.88

= 6.7334g